a the values of or-g-2 ✓-25 is Care) ? b if (3+j2)+(2+J b)=5.54 b=?

Advanced engineering mathimatics
Answer #1

از - آمل (ه || 2 = 6 - 5 - 4 -15 2- 64 -1 - 275 2 واز-8 از = (ن)( ک) 2 - تا(ع) 6 - 2 اذ-5 = (ان+2) + (ان+3) 8 (ط ا ذ - 5 - (b

If you have any doubt please make a comment. Thank you